Meija Isetan Corp. of Japan has two regional divisions with headquarters in Osaka and Yokohama. Selected data on the two divisions follow:

Osaka

Yokohama

Sales

$3,000,000

$9,000,000

Net Operating Income

$210,000

$720,000

Average Operating Assets

$1,000,000

$4,000,000

Which one of the statements is false?

Group of answer choices

Yokohama's profit margin is 7%

Osaka 's asset turnover is 3

Yokohama's asset turnover is 2.25

Osaka's return on assets is 21%

Yokohama's return on assets is 18%
Answer #1

The correct answer is :

Yokohama's profit margin is 7%

Notes:

1. Yokohama's profit margin is 7% = False.

Profit Margin = Net Operating Income/ Sales *100

= 720,000/ $ 9,000,000 * 100

= 8%

Hence,Yokohama's profit margin is 8%

2. Assets Turnover = Net Sales / Average Total Assets

Osaka 's asset turnover is 3 = True

Asset Turnover = $ 3000000 / 1000000

Yokohama's asset turnover is 2.25 =True

Yokohama's asset turnover = 9000000 / 4000000

3. return on assets = Net Income / Average Total Assets

Osaka 's  return on assets = 210000 / 1000000

= 21%

Hence True

Yokohama's return on assets = 720000 / 4000000

= 18%

Hence True